OpenStudy (anonymous):

The solutions of the equation X^2 - 6X + d = 0 are both integers. d is a prime number. Find d.

OpenStudy (anonymous):

The solutions of the equation x2 – 6x + d = 0 are both integers. d is a prime number. d = 5 gives x = 1, 5 ANSWER from x^2 -- 6x + d = 0 = x^2 -- 6x + 5 = (x -- 1)(x -- 5)